OpenStudy (aravindg):

If 4a^2+9b^2-c^2+12ab=0 , then family of straight lines ax+by+c=0 is concurrent at?

OpenStudy (aravindg):

@Miracrown

Miracrown (miracrown):

Any ideas how to begin? O.o

OpenStudy (aravindg):

Concept of pair of straight lines? o.O :-[

OpenStudy (aravindg):

2a+3b=+-c

OpenStudy (aravindg):

2a+3b+c=0 -2a-3b+c=0 What do I do then?

Miracrown (miracrown):

Are we given any restrictions on the values of a, b, and c besides the first equation?

OpenStudy (aravindg):

No restrictions.

Miracrown (miracrown):

What about the 12ab term?

OpenStudy (aravindg):

3*2*2=12

OpenStudy (aravindg):

Dang I got the answer :P

OpenStudy (aravindg):

(2,3) or (-2,-3)

Miracrown (miracrown):

ha ha, really ?

OpenStudy (aravindg):

yeah both satisfy the equation so those are the points :|

OpenStudy (anonymous):

(-2,-3)

Miracrown (miracrown):

yup!

OpenStudy (aravindg):

@sourwing 2 points right?(2,3) or (-2,-3)

Miracrown (miracrown):

It seems to imply that there is some point where all lines of the form ax+by+c=0 intersect if a, b, and c satisfy the equation given.

OpenStudy (anonymous):

2a+3b+c = 0 ax + by + c = 0 a(x+2) + b(y+3) + 2c = 0 or 2a+3b-c = 0 ax + by + c = 0 a(x+2) + b(y+3) = 0 either way, both equations of the the lines go through (-2,-3)

Miracrown (miracrown):